Operations Report – Sept 4th

Off season  Operations Report 4 Sept 2019

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Hab water filter is black

Notes on non-nominal systems:  the water filter is black. I don’t know that replacing it now would do anything so we will need to clean the system before any use in Oct.

Generator (hours run): not run

Generator SOC currently 89% at 6 pm

Diesel Reading – 71%

Propane Reading – Shannon’s  20%.  Main tank 41 %

Ethanol Free Gasoline – 8 gallons.

Water (Axillary tank)  – 0 gallons

Water (static tank) – 400 gallons. 150 gallons at Shannon’s

Axillary to Static tank transfer– no

Gallons transferred: none

Water in GreenHab – n/a gallons

Water (loft) – Static to Loft Pump used – no

Water Meter: n/a

Toilet tank emptied: no

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Notes on rovers: all working rovers in town being used and well cared for.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

No

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: zero

Notes on ATVs: nothing to report

HabCar used and why, where? Not used

CrewCar used and why, where? Yes. To town for drinking water.

General notes and comments:  Productive meeting with Ikea (SLC).  We worked on changes to the plans provided by Ikea in Sweden to improve function over form. The plan is to do the upgrades in three stages: lower deck and GreenHab first to be completed during the October work party. A new kitchen and science lab will be installed later in the season.

Summary of internet: nothing to report

Summary of suits and radios: nothing to report

Summary of Hab operations: nothing to report

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome: Dual split unit was installed in the interior but need the pad for the outside unit which did not get ordered. Once that is herewe will contact an Hvac company to add the refrigerant and then the Science Dome will have both ac and heat. This was installed mainly so we can keep the power system cool in the summer as the dome is pretty well insulated.

Summary of RAM operations: nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: continues to be unseasonably warm. I either have a cold or allergies.

Questions, concerns and requests to Mission Support:

Correction from last night’s report: power lost on June 27th, not July.

Operations Report – Sept 3rd

Off season  Operations Report 3 Sept 2019

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: two GreenHab breakers (new problem), on ENG breaker ( known problem, new reaction)

Notes on non-nominal systems: the power system failed on July 27th (based on data from my Dish Hopper). Three breakers in the Hab tripped as noted above and the short caused a stack mode fault that shut down the system. We got the Science Dome power on first by turning off all the breakers in the Hab, the throwing them one at a time. the ENG breaker is the plugs on the outer wall, which have needed replacement but we were waiting for the remodel. Apparently they shorted. The GreenHab issue is a puzzle and we will start by replacing the two breakers with new ones.

Generator (hours run): one hour. Generator has a small diesel leak that will need to be reported when the season starts. Battery needed to be jumped.

Generator SOC currently 85 % at 6 pm with my AC on

Diesel Reading – 71%

Propane Reading – Shannon’s is currently at 20% and will need a refill although we will probably need a new larger one due to the Intern trailer purchase.  Main tank 41 %

Ethanol Free Gasoline – 8 gallons.

Water (Axillary tank)  – 0 gallons

Water (static tank) – 400 gallons, will need cleaning next month, 150 gallons at Shannon’s

Axillary to Static tank transfer– no

Gallons transferred: none

Water in GreenHab – n/a gallons

Water (loft) – Static to Loft Pump used – no

Water Meter: n/a

Toilet tank emptied: no

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: in town

Hours:

Beginning charge:

Ending charge:

Currently charging:

Notes on rovers: all working rovers in town being used and well cared for.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: all were started and run for awhile. The 300 had to be jumped.

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: one

Notes on ATVs: nothing to report

HabCar used and why, where? Yes. Jumped to start, used to jump generator for start

CrewCar used and why, where? Yes. Using it to travel back and forth to town as the road is too rough for my little rental car.

General notes and comments:  the campus is in decent shape, just has that sad abandoned quality it gets when no one is here. Little evidence of rodents, a few droppings in Hab’s lower deck with none anywhere else. I will have a list of things getting repaired before the work party by the end of the week. We also have a new fifth wheel traler coming on campus for interns and grad students. It’s pretty much a match for mine and will be parked along side mine. Thanks to the donor who supported its purchase, we will have a more active intern program this year.

Summary of internet: all three accounts nominal. We changed the router as it wasn’t working to the old Science one. I don’t have my computer so can’t use mine as the WiFi isn’t turned on. One note of interest:  there is a new cell tower in town and I have some limite text and web capabilities.

Summary of suits and radios: suits are still with NorCal. Radios are in bad shape. Most will need to be replaced

Summary of Hab operations: nothing to report except whoever left the rubber snake has freaked all of us out.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: we will start installing the dual split ac/ heat unit in the ScienceDome tomorrow

Summary of RAM operations: nothing to report

Summary of any observatory issues: Peter confirms robotic Observatory is back on line.

Summary of health and safety issues: it’s very hot. With no power last night and 100 degrees in my trailer when I went to bed I slept poorly. Still hot today even with ac.

Questions, concerns and requests to Mission Support:

Tomorrow we will have a visit from two people from SLC Ikea to work on the ongoing redesign project we have with them.